# Hyper compressed fine-tuning of large foundation models with Quantum Inspired Adapters

### v1

@@ 2.1 Transformer Architecture
For an input sequence $X\in\mathbb{R}^{n\times d}$, where $n,d$ are the sequence length and hidden dimension respectively, the self-attention mechanism computes:
$\text{Attn}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$ (1)
The query, key and value matrices, $Q=XW_{Q}$, $K=XW_{K}$, and $V=XW_{V}$ are linear projections of the input $X$ using learnable weight matrices $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$ respectively.
@@
A typical FFN layer involves two trainable weight matrices, $W_{1}\in\mathbb{R}^{d\times d_{F}}$, $W_{2}\in\mathbb{R}^{d_{F}\times d}$, and is defined as:
$\text{FFN}(X)=\sigma(XW_{1}+\boldsymbol{b}_{1})W_{2}+\boldsymbol{b}_{2},$ (2)
where $d_{F}$ is the dimension of the feed-forward layer and $\sigma$ is a non-linear function which we assume to be $\sigma:=\mathsf{ReLU}$.
2.2 Parameter-Efficient Fine-Tuning Methods
In this section, we review some common and powerful PEFT methods, including Low-Rank Adaptation (LoRA) and Orthogonal Fine Tuning (OFT) and its butterfly variation, (BOFT), which fine-tune orthogonal parameterizations of weight matrices.
2.2.1 Low-Rank Adaptation (LoRA)
There are two main families of adapter configurations one may encounter. The first, are additive adapters, of which Low-Rank Adaptation (LoRA) [5] is an example. Assume a pre-trained weight matrix $W^{*}\in\mathbb{R}^{d_{O}\times d_{I}}$ producing output $\boldsymbol{h}\in\mathbb{R}^{d_{O}}$ given input $\boldsymbol{x}\in\mathbb{R}^{d_{I}}$, $\boldsymbol{h}:=W^{*}\boldsymbol{x}$ with input and output dimensions, $d_{I},d_{O}$ respectively. This weight matrix could be the weights of the FFN $W_{1}$, $W_{2}$ as above or the QKV weight matrices $W_{Q},W_{K},W_{V}$ in the attention mechanism. Adapters introduce additional trainable parameters to the pre-trained model without modifying the original weights. This adapter family is additive because the final model (for inference) is of the form:
$W_{\text{adapt}}:=W^{*}+\Delta W$ (3)
where $\Delta W\in\mathbb{R}^{d_{O}\times d_{I}}$ is the trainable adapter. Note, the adapter matrix, once trained $(\Delta W\rightarrow\Delta W^{*})$, can be merged with the original matrix, and as such does not add any inference overhead. The output of the trained, ‘adapted’, layer is then:
$\boldsymbol{h}_{\text{adapt}}^{*}=W^{*}_{\text{adapt}}\boldsymbol{x}=(W^{*}+\Delta W^{*})\boldsymbol{x}$ (4)
Typically, an adapter module consists of a parameterization that leads to significantly fewer resource requirements, while retaining the outer dimensions $d_{O},d_{I}$. LoRA [5] injects low-rank trainable matrices into the transformer layers by decomposing the weight updates into low-rank matrices. Specifically, the adapter for LoRA are defined as:
$\Delta W_{\text{LoRA}}:=\alpha W_{\text{up}}W_{\text{down}},$ (5)
where $W_{\text{up}}\in\mathbb{R}^{d_{O}\times r}$, $W_{\text{down}}\in\mathbb{R}^{r\times d_{I}}$. Assuming, square pre-trained matrices, $d_{O}=d_{I}=:d$ they enforce that $r\ll d$ ($r$ is the rank), and $\alpha$ is a scaling factor. During fine-tuning, only $W_{\text{up}}$ and $W_{\text{down}}$ are updated.
2.2.2 Orthogonal Fine-Tuning (OFT)
Orthogonal Fine-Tuning (OFT) [8] is an alternative approach to parameter-efficient fine-tuning which enforces an orthogonality constraint on the adapter. The authors justify orthogonality as a useful feature in helping preserve the hyperspherical energy i.e. the angular feature difference between neurons  [16] which in turn helps preserve original knowledge of the model. Unlike methods such as LoRA that inject low-rank updates in an additive manner, OFT and its variants introduce multiplicative adapters. In this case, the updated weight matrix is expressed as:
@@ 2.2.2 Orthogonal Fine-Tuning (OFT)
Again, they assume $W^{*}\in\mathbb{R}^{d\times d}$ is the square pre-trained weight matrix and $\Delta W_{\text{OFT}}\in\mathbb{R}^{d\times d}$ is the orthogonal adapter. Specifically, they have $\Delta W_{\text{OFT}}^{\top}\Delta W_{\text{OFT}}=\mathds{1}$. The orthogonality of $\Delta W_{\text{OFT}}$ ensures that the transformation preserves the spectral properties of $W^{*}$, retaining the pre-trained knowledge during fine-tuning. Different parameterizations of $\Delta W_{\text{OFT}}$ are possible - specifically, [8] chooses to employ the Cayley transform. Given a parameterized matrix, $P\in\mathbb{R}^{d\times d}$, the OFT adapter with the Cayley transform is defined as:
@@
2.2.3 Butterfly Orthogonal Fine-Tuning (BOFT)
Butterfly Orthogonal Fine-Tuning (BOFT) [9] extends OFT by introducing an efficient parameterization of the orthogonal matrix using butterfly structures. In BOFT, the orthogonal matrix $\Delta W_{\text{BOFT}}\in\mathbb{R}^{d\times d}$ is constructed as a product of $m$ sparse orthogonal matrices derived from ‘butterfly’ structures:
@@ 2.2.3 Butterfly Orthogonal Fine-Tuning (BOFT)
where each $\widetilde{B}_{(i)}\in\mathbb{R}^{d\times d}$ is a butterfly factor - a sparse orthogonal matrix that efficiently captures global interactions within the data. These butterfly factors are recursively defined and constructed to ensure orthogonality. The butterfly structure originates from the Cooley-Tukey algorithm for the Fast Fourier Transform, known for its efficient information exchange properties. In BOFT, the butterfly factors are built using small orthogonal blocks that are combined to form larger orthogonal matrices. Specifically, each butterfly factor $\widetilde{B}_{(i)}$ is defined as, $\widetilde{B}_{(i)}=\operatorname{Permute}\left(\operatorname{diag}\left(\Delta W_{\text{BF},1}^{(i)},\Delta W_{\text{BF},2}^{(i)},\dots,\Delta W_{\text{BF},k}^{(i)}\right)\right)$, where $\Delta W_{\text{BF},j}^{(i)}\in\mathbb{R}^{b\times b}$ are small orthogonal matrices parameterized via the Cayley transform (7), $k:=\nicefrac{{d}}{{b}}$ are the number of blocks at level $i$ and $\operatorname{Permute}(\cdot)$ rearranges the blocks to create the butterfly pattern. They typically take the number of butterfly factors to be $m=\log_{b}d$ where $b$ is the block size, and $b\geq 2$. The number of parameters required is $N^{\text{BOFT}}_{P}=\frac{1}{2}md(b-1)=\frac{1}{2}(b-1)d\log_{b}d$ [9]. When $b=2$, the parameter count becomes $N^{\text{BOFT}}_{P}=\mathcal{O}(d\log d)$, compared to the $N^{\text{OFT}}_{P}=\mathcal{O}(d^{2})$ parameters required for a full orthogonal matrix in OFT.
3 Quantum-Inspired Adapters
In this section, we introduce our Quantum-Inspired Adapters, which leverage Hamming-weight preserving quantum circuits prevalent in quantum machine learning literature [10, 11, 12, 17, 18, 19]. Hamming-weight preserving circuits are efficiently implementable quantum circuits that encode and transform data in fixed Hamming-weight subspaces. These circuits have been found useful in quantum machine learning for favorable properties such as absence of barren plateaus [12, 20, 21]. They can be classically implemented using compound matrices, that enforce and retain orthogonality in a natural way via decompositions into Givens rotations [11]. Inspired by these principles, we propose to construct adapters using compound matrices up to a certain Hamming-weight $k$. Combining compounding with orthogonality allows us to create novel adapters which are both expressive and parameter-efficient.
3.1 Construction of Quantum-Inspired Adapters
3.1.1 Compound matrices
Given a ‘base’ matrix, $A\in\mathbb{R}^{n\times n}$, the compound matrix, $\mathcal{C}_{k}:=A^{(k)}$, of ‘order’ $k\in\left[n\right]$ is defined as the $\binom{n}{k}\times\binom{n}{k}$ dimensional matrix with entries $A_{IJ}^{(k)}:=\det(A_{IJ})$ such that $I$ and $J$ are all possible row and column subsets. We use $\mathcal{C}_{k}$ as compact notation for our experiments later in the text. The work of [10] demonstrated how the action of these matrices on different Hamming-weight (different orders, $k$) could be efficiently performed using quantum circuits composed of so-called fermionic beam splitter (FBS) quantum gates. We will describe the quantum implementation in further detail later in the text.
However, we say that the compound adapters which serve the basis of our proposal are quantum-inspired because, for a constant Hamming-weight $k=\mathcal{O}(1)$, the action of these layers can be efficiently classically simulated by direct simulation of the subspaces. We will primarily deal with small order (and combinations thereof) compound matrices in this work, though we leave the open possibility of quantum speedups by quantum implementation of compound layers [12] to future work.
3.1.2 Implementation details
Given a pre-trained weight matrix $W^{*}\in\mathbb{R}^{d\times d}$, we aim to construct a quantum adapter $\Delta W_{Q}\in\mathbb{R}^{d\times d}$ such that $W_{\text{adapt}}=\Delta W_{Q}W^{*}$. Note, here was assume the quantum(-inspired) adapter is a real matrix, which is possible to emulate on quantum circuits despite the existence of complex numbers [11, 10], although we do not rule out the extension of into complex adapters/layers in the future - for example when constructing quantum adapters from general initial states and operations.
Now, the Quantum-Inspired Adapter $\Delta W_{Q}$ is constructed as a block-diagonal matrix:
$\Delta W_{Q}=\bigoplus_{i=1}^{r}\Delta W_{Q}^{i},$ (10)
with $\Delta W_{Q}^{i}\in\mathbb{R}^{b\times b},\forall i$. This decomposition, similarly to OFT, introduces a hyperparameter, $b:=\nicefrac{{d}}{{r}}$, to regulate the total number of parameters. Now each block is itself a block-diagonal matrix constructed using (multiple) compound matrices up to a fixed maximum Hamming-weight, $K$. Temporarily, we define:
$\Delta W_{Q}^{i,*}:=\bigoplus_{k=1}^{K}A^{(k)},$ (11)
At this point, the dimensions of the above objects should be noted. For example, taking a pre-trained weight matrix of size $d=1024$, and assume we have a single block in (10), $N=1$ and a maximum Hamming-weight $K=2$. Then, one method to match dimensions is to choose the dimension of the base matrix, $n$, such that $d-(\text{dim}(A^{(1)})+(\text{dim}(A^{(2)}))=d-\binom{n}{1}-\binom{n}{2}=1024-n-\frac{1}{2}n(n-1)>0$ is minimized, but also an integer. In this example, a suitable choice of $n$ could be $44$. One needs to then finally pad the remaining dimensions which we choose to do as an identity matrix of appropriate size. We leave other choices, and other methods of dimension matching (for example minimizing instead $\mathbb{N}\ni\binom{n}{1}+\binom{n}{2}-d>0$ and truncating the hanging dimensions) for future study.
Figure 2: Different possible adapter configurations. The adapter decomposition is determined by the number of blocks ($b$, or equivalently the ‘rank’ $r:=\nicefrac{{d}}{{b}}$), and the number of compounds within each block. Trailing dimensions are padded with an identity matrix, and are not trainable. The figure shows a) $\mathcal{C}_{1}$, $b=4$ blocks, b) $\mathcal{C}_{1}\oplus\mathcal{C}_{2}$, $b=4$ blocks and c) $\mathcal{C}_{1}\oplus\mathcal{C}_{2}\oplus\mathcal{C}_{3}$, $b=2$ blocks. If the base matrix, $A$ and hence $A^{(1)}=:\mathcal{C}_{1}$ for each block is orthogonal, $A^{\top}A=\mathds{1}$, configuration (a) is equivalent to OFT.
Concretely, in general, we will choose the largest $n$ such that $\sum_{k=1}^{K}\binom{n}{k}\leq b$, construct each $A^{(k)},\forall k\in\{1,\dots K\}$ and then build block, $i$ as:
$\Delta W_{Q}^{i}$ $:=\begin{bmatrix}\Delta W_{Q}^{i,*}&0\\ 0&\mathds{1}_{b-d_{\text{comp}}}\end{bmatrix}=\begin{bmatrix}A^{(1)}&0&\cdots&0\\ 0&A^{(2)}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\mathds{1}_{b-d_{\text{comp}}}\end{bmatrix},$ (12)
where $d_{\text{comp}}:=\sum_{k=1}^{K}\binom{n}{k}$. We show some examples of possible configurations in Figure. 2.
Orthogonality:
Compound matrices, $A^{(k)}$, inherit many properties from their base, $A$ including, invertibility, positive definiteness and, importantly for us, unitarity and orthogonality. By constructing adapter blocks $\Delta W_{Q}^{i}$ using orthogonal compounds and padding with identities, orthogonality is preserved and inherited by $\Delta W_{Q}$. We test the importance of orthogonality as a property for our compound adapters later in section  5.1.
Parameter Efficiency:
The number of trainable parameters are directly controlled by the tuple, $(b,K)$, the number of blocks and number of compounds therein. Note that this is because choosing a larger $K$ reduces the possible size of the base matrix which can be compounded, $n$. All trainable adapter parameters are contained within this base matrix. This results in a compact parameterization suitable for large models. However, the compounding operation builds complex interactions between the parameters in higher orders. We show in our results that this is sufficient to gain high quality results with minimal tuning. However, one could investigate the reordering of features before an adapter layer to directly correspond to different Hamming-weight subspaces which may be more interpretable. For example for graph features, the first order compound, $A^{(1)}$, corresponds to an action on graph nodes, while the second order, $A^{(2)}$ could represent edge interactions. We leave such extensions to future work.
Compounding:
One might ask whether the parameter efficiency is simply result of expanding the effect of a small matrix into a combinatorially large space. As such, the expensive compounding procedure (computing determinants) may not be strictly necessary for performant fine-tuning. However, in the below results we test this hypothesis by replacing the determinant computation (true compound) with alternative operations on matrix minors. For instance, instead of constructing $A^{(k)}$ as $A_{IJ}^{(k)}:=\det(A_{IJ})$, we test the following two element-wise on the matrix minors:
$A_{IJ}^{(k)}=\begin{cases}\max(A_{IJ})\qquad&\text{`}\texttt{max}\text{'}\\ \text{avg}(A_{IJ})\qquad&\text{`}\texttt{avg}\text{'}\end{cases}$ (13)
where $I$ and $J$ are both subsets of rows and columns respectively. We find both of these operations perform worse than the determinant, possibly because they do not respect orthogonality for multiplicative adapters. We leave open the possibility that they may be performant alternatives for additive adapters (e.g. LoRA).
As we will demonstrate in the next sections, our adapters share a crucial flexibility with OFT/BOFT i.e. it’s ability to adapt convolution layers natively. Originally, LoRA is only applied to the linear layers of the attention modules due to it’s additive nature however multiplicative adapters can be used natively when finetuning convolution layers. Moreover, our adapters offer interpretability when applied to convolution layers in the sense that each compound matrix can transform the convolution kernel in a spatial manner. We leave a detailed study of the interpretation of compound adapters in vision tasks for future work.
Table 1: Results on the GLUE development set, fine-tuning the pre-trained DeBERTaV3-base model. # Params denotes the number of trainable parameters. Our method is evaluated with the best configuration, $\mathcal{C^{\prime}}=\mathcal{C}_{1}\oplus\mathcal{C}_{2}$, where orthogonality is enforced ($\gamma=0$), parameter sharing across blocks is disabled ($\beta=0$), and the number of blocks is set to $b=3$.
Method # Params MNLI SST-2 CoLA QQP QNLI RTE MRPC STS-B All
Full Fine-tuning 184M 89.90 95.63 69.19 92.40 94.03 83.75 89.46 91.60 88.25
$\text{LoRA}_{r=8}$ [5] 1.33M 90.65 94.95 69.82 91.99 93.87 85.20 89.95 91.60 88.50
$\text{OFT}_{b=16}$ [8] 0.79M 90.33 96.33 73.91 92.10 94.07 87.36 92.16 91.91 89.77
$\text{BOFT}_{m=2,b=8}$ [9] 0.75M 90.25 96.44 72.95 92.10 94.23 88.81 92.40 91.92 89.89
$\text{Compound}_{\mathcal{C}_{1}\oplus\mathcal{C}_{2}}$ 0.03M 87.46 94.26 64.57 89.08 92.53 81.22 87.99 90.16 85.91
4 Experimental Setup
4.1 Data
We evaluate the effectiveness of our Quantum-Inspired Adapters by finetuning large language models on the General Language Understanding Evaluation (GLUE) benchmark [13], and vision foundation models on the Visual Task Adaptation (VTAB) benchmark [14]. This allows us to evaluate the generalization and robustness of our approach in various language understanding and visual task adaptation challenges.
4.2 Model
We utilize the pre-trained DeBERTaV3-base model [22] as the backbone for our natural language experiments. For vision tasks, we employ the pre-trained DINOv2-large model [23] as our backbone.
4.3 Adapter Configurations

### v2

@@ A.1 Transformer Architecture
For an input sequence $X\in\mathbb{R}^{n\times d}$, where $n,d$ are the sequence length and hidden dimension respectively, the self-attention mechanism computes as follows: $\text{Attn}(Q,K,V)=\text{softmax}\left(\nicefrac{{QK^{\top}}}{{\sqrt{d}}}\right)V$, where the query, key and value matrices, $Q=XW_{Q}$, $K=XW_{K}$, and $V=XW_{V}$ are linear projections of the input $X$ using learnable weight matrices $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d}$ respectively.
@@
A typical FFN layer involves two trainable weight matrices, $W_{1}\in\mathbb{R}^{d\times d_{F}}$, $W_{2}\in\mathbb{R}^{d_{F}\times d}$, and is defined as $\text{FFN}(X)=\sigma(XW_{1}+\mathbf{b}_{1})W_{2}+\mathbf{b}_{2}$, where $d_{F}$ is the dimension of the feed-forward layer and $\sigma$ is a non-linear function which we assume to be $\sigma(\cdot):=\mathsf{ReLU}(\cdot)$.
A.2 Orthogonal Finetuning (OFT)
Orthogonal Finetuning (OFT) [Qiu et al., 2023] is an alternative approach to parameter-efficient finetuning which enforces an orthogonality constraint on the adapter. The authors justify orthogonality as a useful feature in helping preserve the hyperspherical energy i.e. the angular feature difference between neurons  [Liu et al., 2018] which in turn helps preserve original knowledge of the model. Unlike methods such as LoRA that inject low-rank updates in an additive manner, OFT and its variants introduce multiplicative adapters. In this case, the updated weight matrix is expressed as:
@@ A.2 Orthogonal Finetuning (OFT)
Again, OFT assumes $W^{*}\in\mathbb{R}^{d\times d}$ is a square pre-trained weight matrix and $\Delta W_{\text{OFT}}\in\mathbb{R}^{d\times d}$ is the orthogonal adapter, where we have $\Delta W_{\text{OFT}}^{\top}\Delta W_{\text{OFT}}=\mathds{1}$. The orthogonality of $\Delta W_{\text{OFT}}$ ensures that the transformation preserves the spectral properties of $W^{*}$, retaining the pre-trained knowledge during finetuning. Different parameterizations of $\Delta W_{\text{OFT}}$ are possible - specifically, [Qiu et al., 2023] chooses to employ the Cayley transform. Given a parameterized matrix, $P\in\mathbb{R}^{d\times d}$, the OFT adapter with the Cayley transform is defined as:
@@
A.2.1 Butterfly Orthogonal Fine-Tuning (BOFT)
As discussed briefly in the main text, Butterfly Orthogonal Fine-Tuning (BOFT) [Liu et al., 2023] extends OFT by introducing an efficient parameterization of the orthogonal matrix using butterfly structures. In BOFT, the orthogonal matrix $\Delta W_{\text{BOFT}}\in\mathbb{R}^{d\times d}$ is constructed as a product of $m$ sparse orthogonal matrices derived from ‘butterfly’ structures:
@@ A.2.1 Butterfly Orthogonal Fine-Tuning (BOFT)
where each $\widetilde{B}_{(i)}\in\mathbb{R}^{d\times d}$ is a butterfly factor - a sparse orthogonal matrix that efficiently captures global interactions within the data. These butterfly factors are recursively defined and constructed to ensure orthogonality. The butterfly structure originates from the Cooley-Tukey algorithm for the Fast Fourier Transform, known for its efficient information exchange properties. In BOFT, the butterfly factors are built using small orthogonal blocks that are combined to form larger orthogonal matrices. Specifically, each butterfly factor $\widetilde{B}_{(i)}$ is defined as, $\widetilde{B}_{(i)}=\operatorname{Permute}\left(\operatorname{diag}\left(\Delta W_{\text{BF},1}^{(i)},\Delta W_{\text{BF},2}^{(i)},\dots,\Delta W_{\text{BF},k}^{(i)}\right)\right)$, where $\Delta W_{\text{BF},j}^{(i)}\in\mathbb{R}^{b\times b}$ are small orthogonal matrices parameterized via the Cayley transform (4), $k:=\nicefrac{{d}}{{b}}$ are the number of blocks at level $i$ and $\operatorname{Permute}(\cdot)$ rearranges the blocks to create the butterfly pattern. They typically take the number of butterfly factors to be $m=\log_{b}d$ where $b$ is the block size, and $b\geq 2$. The number of parameters required is $N^{\text{BOFT}}_{P}=\frac{1}{2}md(b-1)=\frac{1}{2}(b-1)d\log_{b}d$ [Liu et al., 2023]. When $b=2$, the parameter count becomes $N^{\text{BOFT}}_{P}=\mathcal{O}(d\log d)$, compared to the $N^{\text{OFT}}_{P}=\mathcal{O}(d^{2})$ parameters required for a full orthogonal matrix in OFT.
A.2.2 QuanTA
Here, we give some extended background on Quantum-informed Tensor Adaptation (QuanTA) [Chen et al., 2024], an alternative Quantum-Inspired Adapter recently proposed.
[…]
Tensor networks are decompositions of tensors, i.e. the above QuanTA adapter, $\Delta W_{\text{QuanTA}}\in\mathbb{R}^{d_{1},d_{2},\dots,d_{N},d_{1},d_{2},\dots,d_{N}}$ as a product of smaller tensors usually operating over fewer axes, e.g. three dimensional tensors, $\mathcal{T}\in\mathbb{R}^{d_{r},d_{s},d_{t}}$. The connected graph of $M$ of such tensors is called a tensor network. Tensor networks themselves have found use in machine learning applications for many years, with promising properties for developing and compressing machine learning models [Stoudenmire and Schwab, 2016, Novikov et al., 2015, Tomut et al., 2024].
The full adapter is then constructed by contracting the network over all “virtual” or bond dimensions, and reshaping the “physical” dimensions (i.e. $\{d_{i}\}_{i=1}^{N},\{d_{j}\}_{j=1}^{N}$) back to $d\times d$ for no-overhead inference. As given in [Chen et al., 2024], an $M=3$ tensor example is:
$\Delta W_{\text{QuanTA}}:=\mathcal{T},\mathcal{T}_{i;j}=\mathcal{T}_{i_{1},i_{2},i_{3};j_{1},j_{2},j_{3}}=\sum_{k_{1},k_{2}}\mathcal{T}^{1}_{i_{1},i_{2};k_{1},k_{2}}\sum_{k_{3}}\mathcal{T}^{2}_{k_{1},i_{3};j_{1},k_{3}}\mathcal{T}^{3}_{k_{2},k_{3};j_{2},j_{3}}$ (7)
In the above Eq. (7), each of $\mathcal{T}^{1},\mathcal{T}^{2},\mathcal{T}^{3}$ are 4 index tensors. Here, $\mathcal{T}^{1}/\mathcal{T}^{2}$ carries two/one physical input dimensions, $(i_{1},i_{2})$$i_{3}$ respectively while $\mathcal{T}^{2}$ and $\mathcal{T}^{3}$ carry one/two physical output dimensions, $j_{1}$$(j_{2},j_{3})$ respectively. All other dimensions $(k_{1},k_{2},k_{3})$ are virtual/bond dimensions. Assuming the physical dimensions are fixed, the complexity of dealing with a tensor network contraction (multiplying over bond dimensions) is determined by the dimensions of the bond indices. This also directly regulates the number of trainable parameters within the model/adapter.
Quantum circuit implementation:
Finally, if one wished to translate QuanTA tensors for further quantum-native finetuning (as we discuss in Appendix D) the means of doing so in general is still an open research question. Specifically, quantum computers require unitary operations, and at no stage in training will the tensors in QuanTA have unitarity enforced. Therefore, each of $\mathcal{T}^{1},\mathcal{T}^{2},\mathcal{T}^{3}$ will need to be canonicalised. The canonicalisation procedure makes each tensor an isometry via singular value decomposition through the network. The canonicalisation procedure also enables truncation of the network by clipping singular values. However if the resulting tensors are not square, they will need to be suitably constructed into a full unitary by some method.
Finally, assuming the tensors are not simply two-axes operators (two input and two output qubits), the resulting unitaries need to be compiled to the available gatesets of the quantum computer. One of the most efficient general purpose exact compilation schemes is via the Quantum Shannon Decomposition (QSD) which recursively compiles unitaries into smaller and smaller sub-blocks via de-multiplexing [Shende et al., 2006]. The QSD requires $\nicefrac{{23}}{{48}}\times 4^{n}-\nicefrac{{3}}{{2}}\times 2^{n}+\nicefrac{{4}}{{3}}$ CNOT gates to compile a general $2^{n}\times 2^{n}$ unitary over $n$ qubits, which is exponential in $n$.
Appendix B Technical Proofs
Here, we give the proofs of the Lemmata from the main text.
Lemma (Orthogonality preservation of compound matrices (Lemma 1 repeated)).
If a base matrix, $A\in\mathbb{R}^{n\times n}$ is orthogonal, then all compound matrices, $A^{(k)}$ with $k\in[n]$, for are orthogonal. Furthermore, this orthogonality is preserved during finetuning if we maintain the orthogonality of the base matrix.
Proof.
Let $A\in\mathbb{R}^{n\times n}$ be an orthogonal matrix, i.e., $A^{\top}A=AA^{\top}=\mathds{1}_{n}$. For any $k\in[n]$, the $k$-th compound matrix $A^{(k)}$ has entries $A^{(k)}_{IJ}:=\det(A_{IJ})$ where $I$ and $J$ are $k$-element subsets of $[n]$. Now, to show that $A^{(k)}$ is orthogonal, we need to prove $(A^{(k)})^{T}A^{(k)}=I_{\binom{n}{k}}$.
Consider the $(I,J)$-entry of $(A^{(k)})^{T}A^{(k)}$:
$(A^{(k)}_{IJ})^{\top}A^{(k)}_{IJ}=\sum_{K}A^{(k)}_{KI}\cdot A^{(k)}_{KJ}=\sum_{K}\det(A_{KI})\cdot\det(A_{KJ})$ (8)
By the Cauchy-Binet formula, Eq. 8 equals $\det((A^{\top}A)_{IJ})$. Since $A$ is orthogonal, we have $A^{\top}A=\mathds{1}_{n}$, so:
$\det((A^{\top}A)_{IJ})=\det((I_{n})_{IJ})=\begin{cases}1\qquad\text{if }I=J\\ 0\qquad\text{if }I\neq J\end{cases}$ (9)
Therefore, $(A^{(k)})^{\top}A^{(k)}=\mathds{1}_{\binom{n}{k}}$, proving that $A^{(k)}$ is orthogonal.
To maintain orthogonality during finetuning, we employ the Cayley parameterization as follows. We parameterize the base matrix $A$ using the Cayley transform: $A=(I+Q)(I-Q)^{-1}$ where $Q$ is a skew-symmetric matrix ($Q=-Q^{\top}$). During finetuning, we update only the entries of $Q$ (maintaining its skew-symmetry), which automatically ensures that $A$ remains orthogonal with determinant $1$ (i.e., $A\in\text{SO}(n)$). The compound matrices $A^{(k)}$ are then computed directly from this orthogonal base matrix.
Alternatively, to preserve orthogonality during training, one could employ the quantum strategy of [Landman et al., 2022] described in Appendix D where the orthogonal/compound matrix is trained using its parameterization with Reconfigurable or Fermionic Beam Splitter RBS/FBS quantum gates.
Here we provide a concrete example of how the dimensions of the QuIC adapter components are chosen to match the dimensionality of a pre-trained model’s weight matrix. The primary constraint is that the sum of the dimensions of the compound matrices, $d_{\text{comp}}=\sum_{k=1}^{K}\binom{n}{k}$, must be less than or equal to the block size, $b$. The base matrix dimension, $n$, is typically chosen to maximize this sum without exceeding $b$.
For example, consider a pre-trained weight matrix of size $d=1024$, which we will adapt with a single block ($N=1$, so $b=1024$). If we choose a maximum Hamming-weight of $K=2$, we need to find an integer $n$ such that $\binom{n}{1}+\binom{n}{2}\leq 1024$. To maximize parameterization, we want the largest such $n$. The expression is $n+\frac{n(n-1)}{2}\leq 1024$. A suitable choice is $n=44$, which gives $d_{\text{comp}}=44+\binom{44}{2}=44+946=990$.
The identity matrix $\mathds{1}_{b-d_{\text{comp}}}$ is then added to pad the remaining 1024990=341024-990=34 dimensions. With this example, the matrices in the block defined in Eq. 2 have the following dimensions: $A^{(1)}\in\mathbb{R}^{44\times 44}$, $A^{(2)}\in\mathbb{R}^{946\times 946}$, and the padding identity is $\mathds{1}_{34}\in\mathbb{R}^{34\times 34}$.
Alternatively, if we wished to maximize the number of compound orders for the same block size ($b=1024$), we could choose $n=11$ and $K=5$. This would yield compound matrices $A^{(1)}\in\mathbb{R}^{11\times 11}$, $A^{(2)}\in\mathbb{R}^{55\times 55}$, $A^{(3)}\in\mathbb{R}^{165\times 165}$, $A^{(4)}\in\mathbb{R}^{330\times 330}$, and $A^{(5)}\in\mathbb{R}^{462\times 462}$. The total dimension would be $d_{\text{comp}}=1023$, requiring only a single padding dimension ($\mathds{1}_{1}=1$).
Lemma (Computational Complexity of QuIC Adapters (Lemma 3 repeated)).
Let a QuIC adapter $\Delta W_{Q}$ be defined for a layer of dimension $d$ with $N$ blocks, derived from a base matrix of size $n\times n$ and max compound order $K$.
  1. 1.
    The complexity of the forward pass (applying $\Delta W_{Q}$) is $\mathcal{O}(d^{2}/N)$.
  2. 2.
    The construction of $\Delta W_{Q}$ is a one-time cost, polynomial in $n$ for constant $K$. If parameters are shared, this cost is incurred once per layer.
Proof.
We prove the two parts of the lemma separately.
1. Forward Pass Complexity: The QuIC adapter $\Delta W_{Q}$ has a block-diagonal structure with $N$ blocks, each of size $b\times b$ where $b=d/N$. Applying this adapter to a vector involves $N$ independent multiplications with these smaller blocks. The cost for one block is $\mathcal{O}(b^{2})$. The total cost is therefore:
$N\times\mathcal{O}(b^{2})=N\times\mathcal{O}\left(\left(\frac{d}{N}\right)^{2}\right)=N\times\mathcal{O}\left(\frac{d^{2}}{N^{2}}\right)=\mathcal{O}\left(\frac{d^{2}}{N}\right).$
2. Construction Complexity: The construction of $\Delta W_{Q}$ from the base matrix $A\in\mathbb{R}^{n\times n}$ is dominated by generating the compound matrices $\{A^{(k)}\}_{k=1}^{K}$. To construct the $k$-th compound matrix, $A^{(k)}$, we compute the determinant of all $\binom{n}{k}\times\binom{n}{k}$ minors of size $k\times k$. The cost of a single $k\times k$ determinant is $\mathcal{O}(k^{3})$. Thus, the total cost to construct $A^{(k)}$ is $\mathcal{O}(\binom{n}{k}^{2}\cdot k^{3})$.
The total construction cost sums over all compound orders up to $K$:
$\text{Cost}=\sum_{k=1}^{K}\mathcal{O}\left(\binom{n}{k}^{2}\cdot k^{3}\right).$
For a small, constant maximum order $K$, the complexity is dominated by the largest binomial coefficient term, where $\binom{n}{K}=\mathcal{O}(n^{K})$. The total complexity is therefore $\mathcal{O}(n^{2K})$, which is polynomial in the base matrix size $n$. This construction cost is incurred only once per layer if parameters are shared across blocks, as the resulting matrices can be cached. ∎
Appendix C Adapter Configurations (Extended)